OpenStudy (anonymous):

Graph the data in the table. Draw a line of fit and use any two data points on it to find the slope of the line. Cat Years 1 3 4 6 7 Human Years 13 33 43 63 73 (Points : 1) 20 10 5

OpenStudy (leader):

slope = y2-y1/x2-x1

OpenStudy (leader):

slope=33-13/3-1 slope=20/2=10

OpenStudy (leader):

So slope equals to 10

OpenStudy (leader):

@lee_actress